Your perfectly competitive firm produces 1 million units of output. Your price is $40 per unit and your average total cost is $30 per unit. Your economic profit equals ( in millions)
Answer #1

Economic profit=Total revenue-Total cost

Total Revenue=Price*Quantity=40*1=40 million

Total cost=Average cost*Quantity=30*1=30 million

profit=TR-TC=40-30=10million

Economic profit=10 million